An air column in a pipe which is closed at one end will be in resonance with a vibrating tuning fork of frequency 264HZ if the length of column in cm -
velocity of sound 330m/s

Grade:12th pass

1 Answers

Vikas TU
7 years ago
 
For closed end pipe the Fundamental frequecny tuning fork is given as:
F = v/2l
264 = 330/2l
2l = 330/264
l = 165/264 = 0.625 meter.